# Unleashing HyDRa: Hybrid Fusion, Depth Consistency and Radar for Unified 3D Perception

### v2

@@ 3.2 Height Association Transformer
Table 2: 3D Object Detection on the nuScenes [2] test set. By default, V2-99 [54, 26] is pre-trained for depth estimation, whereas ConvNext-B [44] is initialized from ImageNet-22K [6]. The larger ConvNext Base backbone typically leads to better performance [33].
3.3 Radar-Weighted Depth Consistency
@@
In our method we tackle two challenges of spatial misalignments and projection inconsistencies when integrating the depth information from multiple sensors. CRN [23] tries to compensate for spatial misalignments with a global receptive field with their Multi-modal Deformable Cross Attention (MDCA) module, which solely operates within the BEV. Instead, we propose concatenating the initial set of lifted and splatted BEV features with the radar pillar channels and refining them with our radar-weighted back-projection pipeline. Extending the back-projection-as-refinement [36] to the multi-modal domain, we enforce the consistency between the two projection spaces on both sides of the view disparity [85]. We upgrade the full cycle of Depth Consistency three-fold to enhance the projection quality, leveraging the synergy of our HAT module and a lightweight radar guidance network (RGN). With a small $3\times 3$ convolution followed by a sigmoid, the RGN encodes the radar-only BEV features to additional attention weights $r$ (cf. Fig. 3).
@@
$\text{SCA}(Q_{x,y},F)\!=\!\sum\limits_{i=1}^{{N_{\text{c}}}}\sum\limits_{j=1}^{{N_{\text{ref}}}}\mathcal{F}_{d}(Q_{x,y},\mathcal{P}_{i}(x,y,z_{j}),F_{i})\cdot\bm{w_{c}^{ij}}\cdot\bm{r_{x,y}},$
@@
with $\mathcal{F}_{d}$ denoting the deformable attention, dynamically sampling features around the projection point $\mathcal{P}_{i}(x,y,z_{j})$ on the image feature map $F_{i}$, weighted by depth consistency $w_{c}$ and radar attention $r$. Equation 3 highlights where the complementary hybrid fusion approach is beneficial for a more robust depth prediction. Furthermore, we redesigned the ordering of the modules by moving the backprojection after the temporal fusion to capitalize on partially occluded objects indicated not only by radar reflections but also propagated history features.
Methods Input Backbone Image Size mIoU $\uparrow$ Bus Car Ped. Truck MC
MonoScene [3] C R101 $900\times 1600$ 6.1 4.9 9.4 3.0 7.2 4.0
TPVFormer [20] C R101 $900\times 1600$ 27.8 40.8 45.9 18.9 34.2 20.0
BEVFormerOcc [34] C R101 $900\times 1600$ 26.9 40.4 42.4 21.8 30.7 23.9
CTF-Occ [61] C R101 $900\times 1600$ 28.5 38.3 42.2 22.7 31.1 24.5
SparseOcc [40] C R50 $256\times 704$ 30.9 32.9 43.3 23.4 29.3 23.8
BEVStereo-OCC [18] C R50 $256\times 704$ 36.1 42.1 49.6 21.5 37.1 17.4
OctreeOcc [48] C R50 $256\times 704$ 37.4 - - - - -
FlashOcc (M5) [81] C R50 $256\times 704$ 39.0 41.6 50.5 23.8 37.4 23.0
RenderOcc [53] C R50 $256\times 704$ 39.7 - - - - -
FB-OCC [36, 35] C R50 $256\times 704$ 40.7 45.3 50.5 28.2 37.9 29.0
HyDRa C+R R50 $256\times 704$ 44.4 52.3 56.3 35.1 44.1 35.9
Table 3: 3D Occupancy Prediction on Occ3D based on nuScenes val set (CVPR2023 Occupancy Challenge [61] [62]). In addition to the mIoU based on 16 different classes, we also highlight the IoU of the five dynamic classes car, bus, pedestrian (ped.), truck and motorcycle (MC). An overview of all classes (including drivable area and vegetation etc.) can be found in the Appendix.
3.4 Down-stream Tasks
Our final BEV representation can be leveraged for multiple 3D perception tasks, generalizing well for 3D object detection, 3D multi-object tracking (MOT), and 3D semantic occupancy prediction. To decode the features into bounding box characteristics (3D location, 3D size, yaw orientation, and 2D velocity vector), we follow standard practice for dense architectures and use a simple anchor-free center-based head [80]. Similar to [23], we utilize the same box-based tracking-by-detection approach of CenterPoint [80], which relies on a velocity-based greedy distance matching.
@@ 3.4 Down-stream Tasks
Our powerful fused BEV features can be directly converted into a rich semantic occupancy output without the need to forward projecting features into the full 3D voxel cube or a 3D-convolution-based BEV encoder. We use a one-layer 1$\times$1 Channel-To-Height convolution [81] to create the final representation for our occupancy head [35]. We simply enlarge the final BEV features $F\in\mathbb{R}^{B\times X\times Y\times C}$ in the channel dimension $C$ to $C^{\prime}=Z\times C$ and reshape the feature map to the new height dimension $Z$, resulting in $F^{\prime}\in\mathbb{R}^{B\times X\times Y\times Z\times C}$. This new unstacked representation is fed into the low-cost occupancy head [35].
4 Experiments
4.1 Dataset and Metrics
We perform extensive experiments and ablations on HyDRa  using the nuScenes dataset [2] and the Occ3D benchmark [61] and challenge our model with other relevant camera- and radar-based methods on three different tasks.
nuScenes The nuScenes dataset [2] is the main large-scale research benchmark for camera-radar based 3D perception. The dataset contains 1000 diverse and complex scenarios from the Boston and Singapore urban environments, containing left- and right-handed traffic and adverse conditions like night and rain. Annotated at 2Hz, the dataset provides 6 RGB cameras, 5 mmWave radar sensors, and 1 LiDAR sensor. nuScenes provides a rich set of metrics to assess the quality of 3D perception algorithms: mean Average Precision [10], mean Average Translation, Scale, Orientation, Velocity, and Attribute Error. The NDS is a weighted average of all other metrics. The AMOTA score [71] ranks trackers on the benchmark and evaluates the overall 3D multi object tracking performance.
Occ3D The nuScenes Occ3D benchmark [61, 62] evaluates the quality of arbitrarily shaped 3D foreground and background objects. The goal is to predict the complete 3D scene geometry around the ego-car, consisting of a voxelized representation for 18 classes. The benchmark reports the mean Intersection over Union (mIoU) over the number of true positive, false positive, and false negative voxel predictions.
4.2 Implementation Details
@@
Referring to the literature, we train our model in three different scaling settings following the best practices of previous work. We set the default backbones to ResNet50 [13] with an input resolution of 256$\times$704, 118 depth categories, and a pillar and BEV grid size of 0.8m, leading to 128$\times$128 BEV space. We leverage the AdamW [47] optimizer with a batch size of 64. Scaling up the model backbones and image resolutions to ResNet101 [13] (512$\times$1408) and V2-99 [26] (640$\times$1600), the BEV space is doubled to a standard of 256$\times$256. For comparison to the latest camera-based methods on the test set, we opt for the smaller and more efficient V2-99 backbone.
@@
We build upon the code base of the BEVDet Series [18] and in particular employ the efficient implementation of BEVPoolV2 [17] as a forward projection view transformer. We train for 20 epochs following the memory-efficient sequential sampling and settings of [55] without CBGS [87]. This reduces the step-to-step sampling diversity but speeds up the training time significantly. Regarding the backward projection module, we follow the configuration of FB-BEV [36]. To trade off the number of parameters, we half the backbone output and depth net middle channels. Our HAT module is implemented by one layer of vanilla self- and cross-attention working on 16-times downsampled height queries. This leads to sequences of 16, 32, and 40 height bins for the three model sizes. For temporal fusion, we follow [36, 39, 67] and use the last eight history frames. Data augmentations are the same as BEVDet [18] and CRN [23], but no test-time augmentations are used. For a fair evaluation, each model is trained only on the respective task and leverages the same task head and supervision signals as the baseline methods.
@@
HyDRa C+R V2-99 58.4 0.950 13996 32950 791
Table 4: 3D Object Tracking on nuScenes test set.
4.3 Main Results
3D Object Detection. In Table 1 we report the results of our HyDRa  model on the nuScenes validation set. With a ResNet50 backbone, we achieve a new state-of-the-art of 58.5 NDS, surpassing the previous best score of 56.0 NDS by CRN [23] by 2.5 NDS. Scaling the model and image resolution to ResNet101 and 512$\times$1408, we maintain the large margin and increase the NDS to 61.7, outperforming all other camera- and radar-based methods on the val split. We want to highlight the impressive improvement in the mean Average Translation Error (mATE) of 0.416, which can be attributed to our model’s robust depth estimation concept. Our proposed HyDRa  surpasses CRN with the more powerful ConvNext-Base[33] backbone by 1.4 NDS due to strengthened translation and velocity estimation, leading to a new state-of-the-art of 64.2 NDS on the nuScenes test set (see Table 2). Utilizing a bigger backbone, e.g., Vit-L [9, 67], our method can surpass voxel-based CenterPoint [80]. Qualitative results are published in the Appendix.
Additionally, we show in Table 5 that HyDRa generalizes on the recent View-of-Delft validation set [52], utilizing a more advanced 4D millimeter-wave radar. HyDRa outperforms the best method by an increase of +4.57 mAP in the entire annotation area and +4.06 mAP in the region of interest (compared to the CRN-inspired LXL model [73]).
3D Multi Object Tracking. We summarize the tracking results in Table 4. Benefitting from accurate object localization and high-quality velocity estimation, the center-based tracking performance of HyDRa  is competitive with the state-of-the-art methods, leading to the best trade-off between false positives (FP), false negatives (FN), and identity switches (IDS) on the nuScenes test set, once more narrowing the gap to LiDAR-based CenterPoint.
Methods Input AP in the Entire Annotated Area (%) AP in the Region of Interest (%)
Car Pedestrian Cyclist mAP Car Pedestrian Cyclist mAP
PointPillar [25] R 37.06 35.04 63.44 45.18 70.15 47.22 85.07 67.48
RadarPillarNet [84] R 39.30 35.10 63.63 46.01 71.65 42.80 83.14 65.86
RCFusion [84] C+R 41.70 38.95 68.31 49.65 71.87 47.50 88.33 69.23
SMURF [79] R 42.31 39.09 71.50 50.97 71.74 50.54 86.87 69.72
LXL [73] C+R 42.33 49.48 77.12 56.31 72.18 58.30 88.31 72.93
HyDRa C+R 52.83 56.57 73.25 60.88 80.65 62.90 87.43 76.99
Table 5: 3D object detection results on View-of-Delft  [52]. The RoI ($(x,y,z)|-4m<x<4m,z<25m$) is the driving corridor close to the ego-vehicle. The IoU thresholds are set to 0.5 for cars and 0.25 for pedestrians and cyclists (official evaluation settings).
3D Occupancy Prediction.
Dense architectures with strong depth prediction excel in occupancy prediction [61, 62]. Our BEV-based architecture showcases superior performance on the Occ3D benchmark, outperforming all other camera-based methods by a large margin of 3.7 mIoU (see Table 3). It works remarkably well on challenging dynamic objects as visualized in Figure Fig. 4. Complementary camera-radar fusion is the key to unlocking the potential of vision-centric 3D sensing. While sparse representations perform well in detection (cf. SparseBEV [39] in Table 2), they underperform in occupancy prediction (cf. Sparseocc [40] in Table 3). This emphasizes the importance of a dense representation for full 3D scene understanding.
4.4 Ablation Studies
@@
CRN [23] C+R 48.9 39.7 0.616 0.361
RDC C+R 52.7 40.5 0.534 0.261
HyDRa C+R 53.6 41.2 0.493 0.257
Table 6: Ablation of fusion strategies: RF - Baseline Range-Fusion of projected radar depth maps; CRN with occupancy-based Radar View Transformation (RVT); HyDRa  with Radar-weighted Depth Consistency (RDC) and Height Association Transformer (HAT) is our full model. We train and evaluate in the front-view setting.
To validate the individual contribution of each of our proposed designs and components, we conducted detailed ablation studies. The ablation studies and baseline training of Tab. 6, Tab. 7, and Tab. 8 are performed in the front-view setting. Input for the model are the three forward-facing cameras and the corresponding radar sensors, which reduces the data and computation size by a factor of two. The ablations are conducted on the nuScenes validation set.
@@
We show in Table 6 that this unified approach already outperforms the RVT of CRN. We indicate, that the velocity estimation is also benefitting from a perspective fusion, compared to pure radar-occupancy-guided transformation of the semantic features. Processing the mostly sparse stripe of radar points with 2D convolutions is neither effective nor efficient. To overcome the limitations of the naive baseline, we propose a more sophisticated approach, the HAT module, which is a novel way of fusing radar and camera features in the perspective view and is capable of significantly improving the depth estimation of camera-based models. Compared to related work, we do not have to rely on heuristic height extensions [46, 51] or neglect it at all [12] by collapsing image features [23].
Effect of Radar Weighted Depth Consistency. Radar-informed depth consistency is a crucial concept of our architecture, showing significant and consistent improvements over the baseline architecture of CRN. Especially adding the HAT module into this paradigm leverages synergistic effects, leading to strong metric depth understanding (cf. Tab. 6). Without temporal or depth supervision, as ablated in Table 7, HyDRa  is the superior concept for increased depth and velocity sensing, showing a decent improvement over the camera-only baseline. This could be important for time-critical situations e.g. collision avoidance [78].
Order Matters. In Table 8, we show that the order of the fusion matters. The radar BEV features should be fused before any 4D aggregation and backward projection. Enriching the queries with radar and temporal information, poses more meaningful cues for refinement, without additional computational cost as the modules are working on the same BEV channel dimensions.
@@
HyDRa C+R 45.8 34.7 0.569 0.482
Table 7: Comparison of BEVDepth, CRN and HyDRa without temporal fusion or depth supervision. We train and evaluate in the front-view setting.
@@
A BP $\rightarrow$ 4D $\rightarrow$ RF C+R 52.0 41.1 0.551 0.261
B RF $\rightarrow$ BP $\rightarrow$ 4D C+R 52.7 41.2 0.527 0.281
C RF $\rightarrow$ 4D $\rightarrow$ BP C+R 53.6 41.2 0.493 0.257
Table 8: Ablation of component order (BP: Backprojection, 4D: Temporal Fusion, RF: Radar BEV-Fusion). Radar Attention for Backprojection is not available in setting A. This reports our full model, including HAT and RDC. We train and evaluate in the front-view setting.
4.5 Limitations
In future endeavors, we want to explore the potential of object-level motion modeling [67] within our dense paradigm, as the ego-motion compensation [16] does not account for highly dynamic objects. Moreover, the orientation poses a bottleneck as sparse radar points only give a rough estimate of the object’s location. Therefore, 4D imaging radars pose an interesting research opportunity [52, 7].
5 Conclusion

### v3

@@ III-B Height Association Transformer
TABLE II: 3D Object Detection on the nuScenes test set. The ConvNext backbone typically leads to better performance [6].
[…]
KPConvPillars [51] R Pillars 13.9 4.9 0.823
CenterFusion [14] C+R DLA34 44.9 32.6 0.631
MVFusion [40] C+R V2-99 51.7 45.3 0.569
CRAFT [39] C+R DLA34 52.3 41.1 0.467
BEVDepth [6] C ConvNeXt-B 60.9 52.0 0.445
CRN [13] C+R ConvNeXt-B 62.4 57.5 0.416
StreamPETR [7] C V2-99 63.6 55.0 0.479
SparseBEV [49] C V2-99 63.6 55.6 0.485
[…]
III-C Radar-Weighted Depth Consistency
@@
In our method we tackle two challenges of spatial misalignments and projection inconsistencies when integrating the depth information from multiple sensors. Previous approaches [13] compensate for spatial misalignment with a global receptive field within the BEV. Instead, we propose concatenating the initial set of lifted and splatted BEV features with the radar pillar channels and refining them with our radar-weighted back-projection module. Extending the back-projection-as-refinement [37] to the multi-modal domain, we enforce the consistency between the two projection spaces on both sides of the view disparity. We upgrade the full cycle of Depth Consistency three-fold to enhance the projection quality, leveraging the synergy of our HAT module and a lightweight radar guidance network (RGN). With a $3\times 3$ convolution followed by a sigmoid, the RGN encodes the radar-only BEV features to additional attention weights $r$ (cf. Fig. 4).
@@
$\text{SCA}(Q_{x,y},F)\!=\!\sum\limits_{i=1}^{{N_{\text{c}}}}\sum\limits_{j=1}^{{N_{\text{ref}}}}\mathcal{F}_{d}(Q_{x,y},\mathcal{P}_{i}(x,y,z_{j}),F_{i})\cdot\boldsymbol{w_{c}^{ij}}\cdot\boldsymbol{r_{x,y}},$
@@
with $\mathcal{F}_{d}$ denoting the deformable attention, dynamically sampling features around the projection point $\mathcal{P}_{i}(x,y,z_{j})$ on the image feature map $F_{i}$, weighted by depth consistency $w_{c}$ and radar attention $r$. Equation 3 highlights where the complementary hybrid fusion approach is beneficial for a more robust depth prediction. Furthermore, we moved the backprojection after the temporal fusion to capitalize on partially occluded objects indicated not only by radar reflections but also propagated history features.
III-D Down-stream Tasks
@@
Our final BEV representation can be leveraged for multiple 3D perception tasks, generalizing well for 3D object detection, 3D multi-object tracking (MOT), and 3D semantic occupancy prediction. To decode the features into 3D bounding boxes characteristics, we follow standard practices and use a simple anchor-free center-based head [48]. Similar to [13], we utilize the same box-based tracking-by-detection approach of CenterPoint [48], which relies on a velocity-based greedy distance matching. Our powerful fused BEV features can be directly converted into a rich semantic occupancy output without the need to forward projecting features into the full 3D voxel cube or a 3D-convolution-based BEV encoder. We use a one-layer 1$\times$1 Channel-To-Height convolution [52] to create the final representation for our occupancy head [53]. We simply enlarge the final BEV features $F\in\mathbb{R}^{B\times X\times Y\times C}$ in the channel dimension $C$ to $C^{\prime}=Z\times C$ and reshape the feature map to the new height dimension $Z$, resulting in $F^{\prime}\in\mathbb{R}^{B\times X\times Y\times Z\times C}$. This new unstacked representation is fed into the low-cost occupancy head [53].
IV Experiments
IV-A Dataset and Metrics
We perform extensive experiments and ablations on nuScenes [54], View-of-Delft [55] and Occ3D [16]. The former two provide a rich set of metrics to assess the quality of 3D perception algorithms: mean Average Precision [56], mean Average Translation, Scale, Orientation, Velocity, and Attribute Error. The NDS is a weighted average of all other metrics. The AMOTA score [57] evaluates the overall 3D multi object tracking performance. On Occ3D the goal is to predict the complete 3D scene geometry, consisting of a voxelized representation for 18 classes. The benchmark reports the mean IoU over the number of true positive, false positive, and false negative voxel predictions.
IV-B Implementation Details
@@
Following the literature, we train our model in three different scaling settings. We set the backbones to ResNet50 [24] with an input resolution of 256$\times$704, 118 depth categories, and a pillar and BEV grid size of 0.8m, leading to 128$\times$128 BEV space. Scaling up the model backbones and image resolutions to ResNet101 [24] (512$\times$1408) and V2-99 [58] (640$\times$1600), the BEV space is doubled to a standard of 256$\times$256. For comparison to the latest camera-based methods on the test set, we opt for the smaller and more efficient V2-99 backbone.
@@
We build upon the code base of the BEVDet Series [31] and train for 20 epochs following the memory-efficient sequential sampling of [9] without CBGS [59]. This reduces the step-to-step sampling diversity but speeds up the training time significantly. To trade off the number of parameters, we half the backbone output and depth net middle channels. Our HAT module is implemented by one layer of self- and cross-attention working on 16-times downsampled height queries. This leads to sequences of 16, 32, and 40 height bins for the three model sizes. For temporal fusion, we follow [37, 49, 7] and use the last eight history frames. Data augmentations are the same as BEVDet [31] and CRN [13], but no test-time augmentations are used. For a fair evaluation, each model is trained only on the respective task and leverages the same task head and supervision signals as the baseline methods.
IV-C Main Results
3D Object Detection. In Table I we report the results of our HyDRa model on the nuScenes validation set. With a ResNet50 backbone, we achieve a new state-of-the-art of 58.5 NDS, surpassing the previous best score of 56.0 NDS by CRN [13] by 2.5 NDS. Scaling the model and image resolution to ResNet101 and 512$\times$1408, we maintain the large margin and increase the NDS to 61.7, outperforming all other camera- and radar-based methods on the val split. We want to highlight the impressive improvement in the mean Average Translation Error (mATE) of 0.416, which can be attributed to our model’s robust depth estimation concept. Our HyDRa surpasses CRN with the more powerful ConvNext-Base[6] backbone by 1.4 NDS due to strengthened translation and velocity estimation, leading to a new state-of-the-art of 64.2 NDS on the nuScenes test set (see Table II).
Additionally, we show in Table III that HyDRa generalizes on the recent View-of-Delft validation set [55], utilizing a more advanced 4D millimeter-wave radar. HyDRa outperforms the best method by an increase of +4.57 mAP in the entire annotation area and +4.06 mAP in the region of interest (compared to the CRN-inspired LXL model [60]).
3D Multi Object Tracking. We summarize the tracking results in Table IV. Benefitting from accurate object localization and high-quality velocity estimation, the center-based tracking performance of HyDRa is competitive with the state-of-the-art methods, leading to the best trade-off between false positives (FP), false negatives (FN), and identity switches (IDS) on the nuScenes test set, once more narrowing the gap to LiDAR-based CenterPoint.
TABLE III: 3D object detection results on View-of-Delft  [55]. The official eval settings differentiates a Region of interest.
Methods Input AP Entire Annotated Area (%) AP Region of Interest (%)
Car Ped. Cyclist mAP Car Ped. Cyclist mAP
PointPillar [45] R 37.1 35.0 63.4 45.2 70.2 47.2 85.1 67.5
RCFusion [61] C+R 41.7 39.0 68.3 49.7 71.9 47.5 88.3 69.2
SMURF [62] R 42.3 39.1 71.5 51.0 71.7 50.5 86.9 69.7
LXL [60] C+R 42.3 49.5 77.1 56.3 72.2 58.3 88.3 72.9
[…]
TABLE IV: 3D Object Tracking on nuScenes test set.
@@ IV-C Main Results
[…]
TABLE V: 3D Occupancy Prediction on Occ3D (CVPR2023 Occupancy Challenge [16, 8]). The mIoU based on 16 different classes. All models use a r50 backbone on a resolution of $256\times 704$.
Methods Input mIoU $\uparrow$ Bus Car Ped. Truck MC
SparseOcc [66] C 30.9 32.9 43.3 23.4 29.3 23.8
BEVStereo-OCC [31] C 36.1 42.1 49.6 21.5 37.1 17.4
FlashOcc (M5) [52] C 39.0 41.6 50.5 23.8 37.4 23.0
RenderOcc [67] C 39.7 - - - - -
FB-OCC [37, 53] C 40.7 45.3 50.5 28.2 37.9 29.0
[…]
3D Occupancy Prediction. Dense architectures with strong depth prediction excel in occupancy prediction [16, 8]. Strengthening the most relevant bottlenecks, HyDRa showcases superior performance on Occ3D, outperforming all other camera-based methods by a large margin of 3.7 mIoU (see Table V). It works remarkably well on challenging dynamic objects as visualized in Fig. 5. Complementary camera-radar fusion is the key to unlocking the full potential of vision-centric 3D sensing.
IV-D Ablation Studies
TABLE VI: Ablation of fusion strategies: Baseline Range-Fusion (RF); CRN with occupancy-based Radar View Transformation (RVT); We train and evaluate in the front-view setting.
@@ IV-D Ablation Studies
CRN incl. RVT [13] C+R 48.9 39.7 0.616 0.361
BEVDepth + RDC C+R 52.7 40.5 0.534 0.261
[…]
To validate the contribution of our proposed designs, we conducted detailed ablation studies on nuScenes (val set). The training of Tab. VI and Tab. VII is performed in the front-view setting. Input for the model are the three forward-facing cameras and the corresponding radar sensors, reducing the data and computation size by a factor of two.
@@
We show in Table VI that this unified approach already outperforms the RVT of CRN. We indicate, that the velocity estimation is also benefitting from a perspective fusion, compared to pure radar-occupancy-guided transformation of the semantic features. However, processing the mostly sparse stripe of radar points with 2D convolutions is neither effective nor efficient. To overcome the limitations of the naive baseline, we propose a more sophisticated approach, the HAT module, which is a novel way of fusing radar and camera features in the perspective view and is capable of significantly improving the depth estimation of camera-based models. Compared to related work, we do not have to rely on heuristic height extensions [41, 14] or collapsing image features [13].
Effect of Radar Weighted Depth Consistency. Radar-informed depth consistency is a crucial concept of our architecture, showing significant and consistent improvements over the baseline architecture of CRN. Especially adding the HAT module into this paradigm leverages synergistic effects, leading to strong metric depth understanding (cf. Tab. VI). Without temporal or depth supervision, as ablated in Table VII, HyDRa is the superior concept for increased depth and velocity sensing, showing a decent improvement over the camera-only baseline. This could be important for time-critical situations e.g. collision avoidance [11].
TABLE VII: Ablation without temporal fusion or depth supervision. We train and evaluate in the front-view setting.
@@
V Conclusion